OpenStudy (anonymous):

Which of the following is the solution set of the given equation? (x - 3) - 2(x + 6) = -5 {-10} {-4} {8}

OpenStudy (anonymous):

Its -4, x - 3 -2x -6 + 5 = 0 -4 -3 + 8 -6 + 5 = 0 -7 + 8 - 6 + 5 =0 -13 + 13 = 0 0 = 0

OpenStudy (anonymous):

Thank you so much!